Your US company has a payment due of €7,000,000 next year. Your CEO wants to implement an options hedge. Should you buy a call or put? Both the call and put strikes are $1.15/€. The call costs $0.02/€ and the put costs $0.04/€. What is the amount of dollars you will pay (factoring in the profit/loss of the option) if the spot rate finishes at $1.20/€ next year?
Answer #1

Payable in one year = Euro 7,000,000

US company will need to buy Euro and sell Dollars

Put option is the right to sell and call option is the right to buy

Hence, he should BUY A CALL on Euro

Option Premium = 7,000,000*0.02 = $140,000

Since the rate on maturity ($1.20/Euro) is higher than the strike price ($1.15/Euro), the option will be exercised.

Hence, total amount paid = 7,000,000*1.15 + 140,000

i.e. $8,190,000

Hence, the answer is

BUY a call

$8,190,000